A point charge of 4µC is 10cm distant from a second point charge of −2 µC. Calculate the magnitude of the force on each charge.


1
Expert's answer
2021-04-19T17:06:10-0400

The force between two point charges can be calculated according to Coulomb's law

"F = k \frac{q_1 q_2}{r^2}",

where "q_1, q_2" are the charges, "r" is the distance between them and "k = 9 \cdot 10^9 \frac{N \cdot m^2}{C^2}"is Coulomb's constant.

Hence, "F = 9 \cdot 10^9 \frac{N \cdot m^2}{C^2} \cdot \frac{4 \cdot 10^{-6} C \cdot 2 \cdot 10^{-6} C}{(0.1 m)^2} \approx 7.2 N"